# Predictive tracking of the NV Center based on external temperature sensors

### v1

@@ Abstract
Introduction.– Over the last two decades, quantum computing has successfully evolved from single-qubit demonstrators to systems with up to hundreds of qubits. There is an intense competition between technologies and research groups have started acquiring small scale hardware. Some of these competing technologies are cold atoms [1], superconducting qubits [2], trapped ions [3], quantum dots [4, 5], defects in solids [6], and photonic systems [7].
@@
The problem has conventionally been approached by using a large tracking window (e.g. $5\mu\text{m}\times 5\mu\text{m}$ parallel to the surface [44]), but this leads to either less accurate position estimates or increased relocation time, without effectively solving the problem. Alternatively, an $8$-point method can track with significantly fewer data points [45, 46]. While current state of the art methods aim to reduce the number of points scanned, they still rely on some form of active relocation measurements, interrupting any quantum computation.
In this work, we determine the NV’s position without requiring active scanning, allowing for continuous operation. We report on the mechanism of placing the optical equipment in thermal isolation leading to the correlation of the NV’s position and resonance frequency with the room’s temperature sensors. By training a simple model from the experimental data, we are able to accurately predict the positions and resonance frequency $\nu_{\mathrm{res}}$ using external temperature measurements as inputs. As we show below, this design and implementation is a simple yet highly effective way to reduce the time spent to periodically track the electron within the NV at room temperature inside a tight window ($1\mu\text{m}\times 0.7\mu\text{m}$) parallel to the surface.
Experiment.– The experimental setup is designed such that the optical table is enclosed in a box of dimensions $45$ cm ×$\times~{}50$ cm×$\times~{}15$ cm (L x W x H). Figure 1 shows the setup. The optics table is located in a small room of dimensions around $2$ m $\times$ $2$ m $\times$ $3$ m that has a weak thermal isolation, i.e. no use of heat exchange equipment. The optics table is liquid cooled and kept at a constant target temperature. A sensor is placed on the optics table to measure the temperature $\text{T}_{1}$. The major sources of heat generation, e.g. pulse generators, are kept outside the room. T2 is the room temperature measured through a single sensor mounted in the room and varies uncontrolled with the temperature outside. Both $\text{T}_{1}$ and $\text{T}_{2}$ are not to be confused with spin relaxation or coherence times and are measured every $20$ seconds.
@@
Refer to caption(a)Refer to caption(c)Refer to caption
@@
Figure 2: (a) The relative change in NV’s location (left axis) and temperature T2 (right axis) plotted against a time period of $10$ days. The values in $X$ and $Z$ directions are negated for a better visualisation of the anticorrelation. There are strong linear (anti/)correlations despite the different scales of $X$, $Y$, and $Z$. The axes are not optimized for superimposition. (b) The correlation matrix between the NV’s coordinates $X$, $Y$, and $Z$, and temperatures T1 and T2 shows the strength of their linear correlations (red) and anticorrelations (blue). The apparent stronger (anti/)correlation of positions with T2 compared to T1 is due to the measurement sensitivities of the sensors. As is also visible in (a), the measurables $-X$, $Y$, and T2 have near perfect correlation. (c) The percentage difference in the crest and trough of the Rabi oscillations obtained for different driving frequencies at constant temperature is plotted in red square points. A Lorentzian fit to the data is shown in dashed black lines. The full width at half maxima (FWHM) is $1.55$ MHz.
We continuously track the change in the location of the NV w.r.t. the temperature over a period of $10$ days. The lowest and highest temperatures measured were $13.08$°C and $13.31$°C for T1, and $15.8$°C and $20.0$°C for T2, respectively. It is worth noting that T1 and T2 have different least counts, namely, $0.01$°C and $0.02$°C, respectively. The point of origin of the NV’s coordinates is chosen arbitrarily. Data collection was interrupted for ten minutes on day three by an unexpected system restart. This caused an unusually large change in $Y$ and $Z$, attributed to an unknown software glitch during the abnormal shutdown. For data normalisation, we added $-0.227\mu$m and $0.05\mu$m to all $Y$ and $Z$ values after the third day, respectively, such that they had identical positions in the last pre- and first post-interruption measurements.
@@
We plot the change in the NV’s location w.r.t. T2 in Fig. 2(a). The left y-axis gives the positions in $-X$, $Y$, and $-Z$ coordinates, respectively, in units of $10^{-6}$ m. The values of $X$ and $Z$ are negated in order to better visualize their anticorrelation w.r.t. T2. We perform a linear interpolation of all the four variables since the temperature data and NV’s location is not collected at exactly the same time but often with a difference of less than few tens of seconds. Fig. 2(a) shows only a finite sample of these interpolated data points for visualization reasons.
@@
To better understand the relationship between the NV’s location and the temperatures, we linearly interpolate $10^{5}$ data points to calculate any correlations between the variables. First, we normalise the data of all $5$ variables such that it ranges between 00 and $1$. Second, we rearrange T2 values in the decreasing order and sort all the other variables according to this order. Doing so allows us to retain the same timestamped values of others variables in reference to T2. Third, we do a linear fit to all the variables to compute the respective slopes. Fourth, we use $1-\frac{2\left|\,|a|-|b|\,\right|}{\left(|a|+|b|\right)}$ or $\frac{2\left|\,|a|-|b|\,\right|}{\left(|a|+|b|\right)}-1$ when slopes $a$ and $b$ have the same signs or opposite signs, respectively, to compute the correlations. Finally, we calculate these correlations for all pairs of variables and show the results in Fig. 2(b).
@@
Refer to caption(a)(b)
Figure 4: (a) Comparison of the actual and predicted $X$ (top), $Y$ (middle), and $Z$ (bottom). The yellow area highlights size of the tracking window. The units of y-axes is meters. (b) The actual $\nu_{\mathrm{res}}$ (grey) in GHz and the predicted $\nu_{\mathrm{res}}$ (orange) as a function of time. Shown in blue are error bars depicting half-contrast off-resonance $\nu$ using FWHM from Fig. 2(c).
@@
We use second order polynomial regression to predict the position coordinates of the NV and the resonance frequency $\nu_{\mathrm{res}}$. We train our models on the data T1 and T2 linearly interpolated to $60$ points per minute across the first two and half days. We use the model
@@
where we estimate the coefficients $\beta_{i,\gamma}$ by minimizing least squares for each ${\gamma}\in\{X,Y,Z,\nu_{\mathrm{res}}\}$ separately. The choice of the model is motivated by the results shown in Fig. 2(a) where some non-linear contributions are apparent to mostly linear correlations. With ordinary least squares, these parameters are optimized to best fit the interpolated observed data and capture both linear, using $\beta_{0,\gamma},\beta_{1,\gamma},\beta_{2,\gamma}$, and non-linear, using $\beta_{3,\gamma},\beta_{4,\gamma},\beta_{5,\gamma}$, contributions. The goal of the modeling is to accurately and precisely predict $\nu_{\mathrm{res}}$ and coordinates $X$, $Y$, and $Z$ given T1 and T2. The accuracy and precision of prediction for the coordinates is relative to the size of the tracking window in which the NV’s position is scanned. We emphasise that a larger tracking window is easily possible but we test our model’s predictions against the same conditions under which it was trained. The accuracy of the prediction for $\nu_{\mathrm{res}}$ is relative to the Rabi contrast shown in Fig. 2(c) and its full width at half maxima (FWHM).
The results for the coordinate predictions are shown in Fig. 4(a). We observe that the variable $X$ moves out of the tracking window after about one day as the temperature starts to increase. Our simple model is able to predict locations for coordinates $X$ and $Y$ very accurately. While our model qualitatively captures the movement of $Z$, it fails to quantitatively predict $Z$ at the same precision as the others. However, the gap between the predicted and actual $Z$ values is well compensated by the size of the tracking window.
@@
In Fig. 4(b), we show the prediction for $\nu_{\mathrm{res}}$ along with error bars calculated from FWHM when the contrast of the Rabi oscillations drops to half its peak value at constant temperature (see Fig. 2(c)). We observe that our model is able to accurately predict the change in $\nu_{\mathrm{res}}$ up to the error limits for slightly more than a week when temperature is left uncontrolled. We emphasize that for continued normal operations of an NV center based quantum computer, it is only required that $\nu_{\mathrm{res}}$ is predicted accurately within a few hours, not days. Our simple model is sufficient for this purpose.
[…]
@@
If T2 varies uncontrolled, as in our experiment, the trained model allows one to predict the NV’s movement without requiring periodic active position measurements. The NV’s position can then be updated, without active scanning, at regular intervals by the model’s predictions based on T1 and T2. This prevents the interruptions of quantum algorithms and allows for continuous operation of the quantum computer.
@@
Conclusion.– We described a simple experimental setup where an NV center in a diamond is placed in weak thermal isolation correlated very well with temperature sensors for measurements taken over ten days. We demonstrated that the resonance frequency is also correlated to the room temperature. By using a simple quadratic model as a function of the temperature of two sensors, training it for a period of two and half days, we were able to accurately and precisely predict the movement of the NV for the next one week. Similarly, our model was able to predict $\nu_{\mathrm{res}}$ to a reasonable accuracy.
@@
Acknowledgement.– We thank Gopalakrishnan Balasubramanian, Priya Balasubramanian, Matthias Gerster, and Florian Frank from XeedQ GmbH for their generous support around the hardware setup. We also thank Philip Döbler for his feedback.

### v2

@@ Abstract
Over the last two decades, quantum computing has evolved from single-qubit demonstrators to systems with up to hundreds of qubits. There is an intense competition between technologies and research groups have started acquiring small scale hardware. Some of these competing technologies are cold atoms [1], superconducting qubits [2], trapped ions [3], quantum dots [4, 5], defects in solids [6], and photonic systems [7].
@@
The problem has conventionally been approached by using a large tracking window (e.g. $5\mu\text{m}\times 5\mu\text{m}$ parallel to the surface [44]), but this leads to either less accurate position estimates or increased relocation time, without effectively solving the problem. Alternatively, multi-point methods can track with significantly fewer data points. The use of 88-points [45, 46], and even 6,4,36-,4-,3-point methods has been demonstrated [47, 48, 49]. In nanodiamond thermometry, more points are expected to be better at tracking but at the expense of less measurements which may lead to decline in precision [47]. In quantum computing, any time spent on tracking the NV is lost for actual computation. While the multi-point methods aim to reduce the number of points scanned, they still rely on some form of active relocation measurements, interrupting any form of computation. Overall, despite a reduction in the track-and-locate time using multi-points, these methods essentially remain in the paradigm of stop-and-fix, where the computation is stopped to fix the tracked coordinates.
In this work, we determine the NV’s position without requiring active scanning, allowing for continuous operation. We report on the mechanism of placing the optical equipment in thermal isolation leading to the correlation of the NV’s position and resonance frequency with the room’s temperature sensors. By training a simple model from the experimental data, we are able to accurately predict the positions and resonance frequency $\nu_{\mathrm{res}}$ using external temperature measurements as inputs. As we show below, this design and implementation is a simple yet highly effective way to reduce the time spent to periodically track the electron within the NV at room temperature inside a tight window ($1\mu\text{m}\times 0.7\mu\text{m}$) parallel to the surface. Our work moves the recalibration into a different paradigm: a simple experimental redesign that enables modelling the drift and predicting it in real-time. The paradigm is moved from stop-and-fix to predict-and-prevent. The location of the NV is predictively tracked and the necessity for scanning is prevented for prolonged periods of time.
The experimental setup is designed such that the optical table is enclosed in a box of dimensions $45$ cm $\times~50$ $\times~15 cm$ (L x W x H). Figure 1 shows the setup. The optics table is located in a small room of dimensions around $2\times~2$ $\times 3 m$ that has a weak thermal isolation, i.e. no use of heat exchange equipment. The optics table is liquid cooled and kept at a constant target temperature. A sensor is placed on the optics table to measure the temperature $\text{T}_{1}$. The major sources of heat generation, e.g. pulse generators, are kept outside in a bigger room which is air cooled. T2 is the room temperature measured through a single sensor mounted in the room and varies uncontrolled with the temperature outside. Both $\text{T}_{1}$ and $\text{T}_{2}$ are not to be confused with spin relaxation or coherence times and are measured every $20$ seconds.
@@
Refer to caption(a)Refer to caption(c)Refer to caption(b)
@@
Figure 2: (a) The relative change in NV’s location (left axis) and temperature T2 (right axis) plotted against a time period of $10$ days. The values in $X$ and $Z$ directions are negated for a better visualisation of the anticorrelation. There are strong linear (anti/)correlations despite the different scales of $X$, $Y$, and $Z$. The axes are not optimized for superimposition. (b) The correlation matrix between the NV’s coordinates $X$, $Y$, and $Z$, and temperatures T1 and T2 shows the strength of their linear correlations (red) and anticorrelations (blue). The apparent stronger (anti/)correlation of positions with T2 compared to T1 is due to the measurement sensitivities of the sensors. As is also visible in (a), the measurables X-X, $Y$, and T2 have near perfect correlation. (c) The percentage difference in the crest and trough of the Rabi oscillations obtained for different driving frequencies at constant temperature is plotted in red square points. A Lorentzian fit to the data is shown in dashed black lines. The full width at half maxima (FWHM) is $1.55$ MHz.
[…]
@@
We plot the change in the NV’s location w.r.t. T2 in Fig. 2(a). The left y-axis gives the positions in X-X, $Y$, and Z-Z coordinates, respectively, in units of $10^{-6}$ m. The values of $X$ and $Z$ are negated in order to better visualize their anticorrelation w.r.t. T2. We perform a linear interpolation of all the four variables since the temperature data and NV’s location is not collected at exactly the same time but often with a difference of less than few tens of seconds. Fig. 2(a) shows only a finite sample of these interpolated data points for visualization reasons.
@@
To better understand the relationship between the NV’s location and the temperatures, we linearly interpolate the data points to calculate any correlations between the variables. First, we normalise the data of all $5$ variables such that it ranges between 0 and $1$. Second, we rearrange T2 values in the decreasing order and sort all the other variables according to this order. Doing so allows us to retain the same timestamped values of others variables in reference to T2. Third, we do a linear fit to all the variables to compute the respective slopes. Fourth, we use $1-\frac{2\left|\,|a|-|b|\,\right|}{\left(|a|+|b|\right)}$ or $\frac{2\left|\,|a|-|b|\,\right|}{\left(|a|+|b|\right)}-1$ when slopes $a$ and $b$ have the same signs or opposite signs, respectively, to compute the correlations. Finally, we calculate these correlations for all pairs of variables and show the results in Fig. 2(b).
@@
Refer to captionRefer to caption(a)(b)
Figure 4: (a) The absolute difference between actual and predicted $X$ (top), $Y$ (middle), and $Z$ (bottom). The units of y-axes is nanometers. (b) The actual $\nu_{\mathrm{res}}$ (grey) in GHz and the predicted $\nu_{\mathrm{res}}$ (orange) as a function of time.
@@
We use second order polynomial regression to predict the position coordinates of the NV and the resonance frequency $\nu_{\mathrm{res}}$. We train our models on the data T1 and T2 suitably linearly interpolated across the first two and half days so that the data points have same timestamps. We use the model
@@
where we estimate the coefficients $\beta_{i,\gamma}$ by minimizing least squares for each ${\gamma}\in\{X,Y,Z,\nu_{\mathrm{res}}\}$ separately. The choice of the model is motivated by the results shown in Fig. 2(a) where some non-linear contributions are apparent to mostly linear correlations. With ordinary least squares, these parameters are optimized to best fit the interpolated observed data and capture both linear, using $\beta_{0,\gamma},\beta_{1,\gamma},\beta_{2,\gamma}$, and non-linear, using $\beta_{3,\gamma},\beta_{4,\gamma},\beta_{5,\gamma}$, contributions. The goal of the modeling is to accurately and precisely predict $\nu_{\mathrm{res}}$ and coordinates $X$, $Y$, and $Z$ given T1 and T2. The accuracy and precision of prediction for the coordinates is related to the size of the tracking window in which the NV’s position is scanned.
The results for the coordinate predictions are shown in Fig. 4(a). The plots show the absolute difference between the actual data and the prediction data of the model. The results show that our model’s accuracy is different in each coordinate. The predictions become less accurate over time, which is to be expected. While the model is able to predict locations for coordinates $X$ and $Y$ rather accurately, it only qualitatively captures the movement of $Z$. The underlying cause requires further investigation. However, the prediction error in $Z$ is still well within the size of the tracking window.
@@
In Fig. 4(b), we show the prediction for $\nu_{\mathrm{res}}$ along with the actual values. We observe that our model is able to predict the change in $\nu_{\mathrm{res}}$ within about one FWHM (see Fig. 2(c)) of Rabi contrast for slightly more than a week when temperature is left uncontrolled. The cause of the frequency drift is most likely the change of distance between the magnet and the diamond due to the thermal expansion/contraction of the optics table on which they are mounted. We emphasize that for continued normal operations of an NV center based quantum computer, it is only required that $\nu_{\mathrm{res}}$ is predicted accurately within a few hours, not days. Our simple model is sufficient for this purpose.
Albeit highly setup-specific, it is interesting to calculate the shift in NV’s position per °C by dividing the maximum change in each coordinate by the maximum change in T2 across the entire data. The position changes by $\Delta X=1.533\mu\text{m/\textdegree C},\Delta Y=0.178\mu\text{m/\textdegree C}$, and $\Delta Z=0.682\mu\text{m/\textdegree C}$. Similarly, $\Delta\nu_{\mathrm{res}}$ = $517$ kHz/°C. Thus, for a temperature change of >0.33>0.33 °C, >1.98>1.98 °C, and >2.21>2.21 °C, the brightest spot of the NV’s point spread function (PSF) will be outside our tracking window in the $X$, $Y$, and $Z$ directions, respectively. Similarly, a temperature change of $1.41$ °C reduces the Rabi contrast by half. These values give us a reference for our experiments to which, if the external temperature T2 is optimally controlled, and everything else remains unchanged, the NV will always be found within the tracking window and reachable at the same resonance frequency with a high contrast.
@@
If T2 varies uncontrolled, as in our experiment, the trained model allows one to predict the NV’s movement without requiring periodic active position measurements. The NV’s position can then be updated, without active scanning, at regular intervals by the model’s predictions based on T1 and T2. This prevents the interruptions of quantum algorithms and allows for continuous operation of the quantum computer. While our model takes into account the temperature fluctuations and makes reasonable predictions for the position coordinates, temperature is not a unique factor influencing the NV’s position. Factors like humidity, diamond strain, chiller flow rate, and table vibrations should be investigated and potentially added to the simple model for more precise predictions.
Besides quantum computing, our method can be useful in nanodiamond thermometry where multi-points methods are currently used. If the nanodiamonds are immobilized, for instance by surface adsorption, polymer embedding, or covalent attachment to a substrate, their Brownian motion can be effectively suppressed, and our approach could in principle be applied in this scenario for real-time estimations of magnetic resonance shifts in biological applications.
@@
We described a simple experimental setup where an NV center in a diamond is placed in weak thermal isolation correlated very well with temperature sensors for measurements taken over ten days. We demonstrated that the resonance frequency is also correlated to the room temperature. By using a simple quadratic model as a function of the temperature of two sensors, training it for a period of two and half days, we were able to accurately and precisely predict the movement of the NV for the next one week. Similarly, our model was able to predict $\nu_{\mathrm{res}}$ to a reasonable accuracy.
@@
We thank Gopalakrishnan Balasubramanian, Priya Balasubramanian, Matthias Gerster, and Florian Frank from XeedQ GmbH for their generous support around the hardware setup. We also thank Philip Döbler for his feedback.